OpenStudy (iwanttogotostanford):

Solve x^2 + 3x + 1 = 0.

OpenStudy (iwanttogotostanford):

@Lord_Frieza

OpenStudy (anonymous):

Im back

OpenStudy (anonymous):

K here are my steps tell me if you dont understand them

OpenStudy (iwanttogotostanford):

ok

OpenStudy (anonymous):

Let's solve your equation step-by-step. x2+3x+1=0 Step 1: Use quadratic formula with a=1, b=3, c=1. x= −b±√b2−4ac 2a x=−(3)±√(3)2−4(1)(1) 2(1) x= −3±√5 2 x= −3 2 +12√5 or x=−32 + −1 2 √5 Answer: x=−32+12√5 or x=−32+−12√5

OpenStudy (iwanttogotostanford):

-32??

OpenStudy (iwanttogotostanford):

that isn't even one of my answers

OpenStudy (anonymous):

omfg sorry

OpenStudy (anonymous):

I fudged up so hard lol

OpenStudy (iwanttogotostanford):

its fine

OpenStudy (anonymous):

can you give me the optional answers plz

OpenStudy (iwanttogotostanford):

x equals the quantity of 3 plus or minus square root 13 all over 2 x equals the quantity of negative 3 plus or minus square root 13 all over 2 x equals the quantity of negative 3 plus or minus square root 5 all over 2 x equals the quantity of 3 plus or minus square root 5 all over 2

OpenStudy (anonymous):

ok lol i typed it wrong heres my asnwer x=−32+1√5 or x=−32+−12√5

OpenStudy (iwanttogotostanford):

thats still not right

OpenStudy (anonymous):

omfg wait one sec replace 32 with -3/2

OpenStudy (anonymous):

i keep forgetting the /

OpenStudy (iwanttogotostanford):

yeah thats right now haha thanks! I get it

OpenStudy (anonymous):

Lol sorry my / key was jammed